OpenStudy (anonymous):

Pam has a complex number with a real part of -2 and an imaginary part of -7. what is the conjugate of pams number? (explain how you got it)

OpenStudy (anonymous):

@thomaster @Luigi0210 @preetha @mathmale

OpenStudy (anonymous):

Help?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@agent0smith @mathslover

OpenStudy (anonymous):

please help me ASAP i really need help @hartnn @Luigi0210 @thomaster @preetha @mathmale @RadEn @mathslover @Compassionate @agent0smith @Ashleyisakitty

OpenStudy (agent0smith):

Conjugate of a + bi is a - bi. Eg. 2 - 4i would have conjugate 2 + 4i real part of -2 and an imaginary part of -7. your number is -2 - 7i

OpenStudy (anonymous):

i dont understand can you try to futher explain so i can understand

OpenStudy (agent0smith):

All you do is change the sign (the + or -) on the imaginary part (with the i).

OpenStudy (anonymous):

so the answer is simply -2-7i

OpenStudy (anonymous):

can you help me with one more problem

OpenStudy (anonymous):

show all the work to simplify. leave your answer in simple radical form: 1.7^ sqrt of x^2 2.(^3 sqrt x^4)^8

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@ParthKohli can you help me?

OpenStudy (anonymous):

please help walter i need help with the 2nd question

Parth (parthkohli):

It's really hard to understand the formatting of the second question...

OpenStudy (anonymous):

sorry ill try to retype it

OpenStudy (anonymous):

\[1.\sqrt[7]{x}^2\] \[2.(\sqrt[3]{x}^4)^8\]

OpenStudy (anonymous):

is that good enough @ParthKohli

Parth (parthkohli):

Ah.

Parth (parthkohli):

Do you know that\[\large \sqrt[a]{x^b} = x^{b/a}\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

so would number 1 be x^7/2

Parth (parthkohli):

OK, now you do. Number one would be\[x^{2/7}\]

OpenStudy (anonymous):

is that the final answer?

Parth (parthkohli):

I believe, yes.

OpenStudy (anonymous):

so number 2.

OpenStudy (anonymous):

would it be (x^4/3)^8

OpenStudy (anonymous):

@ParthKohli ?

Parth (parthkohli):

Yes!

OpenStudy (anonymous):

whats next?

OpenStudy (anonymous):

would i evenly distribute the 8

Parth (parthkohli):

yes, in other words\[(a^m)^n= a^{m \times n}\]

OpenStudy (anonymous):

so x^32/24

OpenStudy (anonymous):

is that the final answer or would i simplify? @ParthKohli

OpenStudy (anonymous):

@ParthKohli

Parth (parthkohli):

That's not right.

Parth (parthkohli):

You only multiply to the numerator.

OpenStudy (anonymous):

so it would be 768? @parthKohli

Parth (parthkohli):

Hmm! I must leave now... sorry!

OpenStudy (anonymous):

whats the answer

OpenStudy (anonymous):

b4 you leave

OpenStudy (anonymous):

@ParthKohli please

OpenStudy (agent0smith):

"so the answer is simply -2-7i" the conjugate of -2-7i is not -2-7i Re-read everything I wrote above.

OpenStudy (anonymous):

oh im sorry its 2+7i my teacher said something about it im sorry for not putting your answer as a best response and i fanned you thnks

OpenStudy (anonymous):

can you actually help me with what my teacher said about my answers@agent0smith

OpenStudy (agent0smith):

I don't know what your teacher said...

OpenStudy (agent0smith):

it's not 2+7i either.

OpenStudy (agent0smith):

Read this carefully: Conjugate of a + bi is a - bi. Eg. 2 - 4i would have conjugate: 2 + 4i your number is -2 - 7i Find its conjugate.

OpenStudy (agent0smith):

Here's another example: -5 + 7i would have conjugate: -5 - 7i

OpenStudy (anonymous):

im sorry my compuyer crashed

OpenStudy (anonymous):

wouldnt it be -2+7i then?

OpenStudy (agent0smith):

Yes

OpenStudy (anonymous):

really awesome do you think ytou could help me with some other problems

OpenStudy (anonymous):

#1A) You are correct that you would move the index to the exponent. However, reverse the fraction. The 2 goes on top, and 7 on the bottom: x^(2/7) #1B) Think of it as: x^(4/3)^(8/1) Then you multiply: 4/3 * 8/1 What do you get for your fraction exponent now?? @agent0smith

OpenStudy (agent0smith):

Post a new question, this is far too long and is unrelated to the original question

OpenStudy (anonymous):

no its just 2 please help i dont want to have to make another post please just help me now and there half don u dont understand certain things in the problems

OpenStudy (agent0smith):

No, i don't want to look for the original problems in this mess. You'd be better off doing them from the start anyway, than trying to figure out where you went wrong.

OpenStudy (anonymous):

but the first one is just \[x^{2/7}\]

OpenStudy (anonymous):

what would i do next divide 7 by 2? @agent0smith

OpenStudy (agent0smith):

You already have the answer

OpenStudy (anonymous):

ok so now 2 wouldnt x^(4/3)^(8/1) i multiple the exponents so it would be x^32/3

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@abb0t

OpenStudy (agent0smith):

yes

OpenStudy (anonymous):

i was correct?

OpenStudy (agent0smith):

yes

OpenStudy (anonymous):

well thnks @agent0smith your help was greatly appreciated i fanned you and will prob get around writing a testimonial